OpenStudy (anonymous):

find value of the integral f [3,1] 1/x^2 dx

OpenStudy (anonymous):

\[\int\limits_{1}^{3} \frac{ 1 }{ x^2 } dx\] \[Deltax=\frac{ 3-1 }{ n }=\frac{ 2 }{ n }\]

OpenStudy (unklerhaukus):

\[\int\limits^3_1\frac1{x^2}\mathrm dx\\ =\int\limits^3_1x^{-2}\mathrm dx\\ =\left.\frac{x^{-2+1}}{-2+1}\right|_1^3\\=\]

OpenStudy (anonymous):

so the value of the integral is 2/3 as f(3)=-1/3 f(1)=-1 so -(-2/3) which = 2/3

OpenStudy (unklerhaukus):

I think you've done this right , because you have the right answer, but im going to show you the steps i would use anyway \[\int\limits^3_1\frac1{x^2}\mathrm dx\\ =\int\limits^3_1x^{-2}\mathrm dx\\ =\left.\frac{x^{-2+1}}{-2+1}\right|_1^3\\ =\left.-{x^{-1}}\right|_1^3\\ =-3^{-1}-(-1^{-1})\\=-\frac13+1\\=1-\frac13\\=\frac23\]